the ratio of sand to cement in a mixture of concrete is 4:5. how much of each ingredient is there in 144kg of this mixture?
could u please give me the way it is done? thank you in advance.

Answer by mananth(16946)

The ratio of cement to sand = 4/5
multiply by a common factor x
4x parts of cement
5x parts of sand.
4x+5x=144
9x=144
x=144/9
x=16
4x/5x = 4*16/5*16= 64/80
cement = 64 kg
Sand = 80 kg
CHECK
64+80 = 144
64/80 = 4/5
m.ananth@hotmail.ca